OpenStudy (anonymous):

34 tafara27 It takes 6hrs longer to cross a channel in a ferryboat when one engine of the boat is usd alone than when a second engine is used alone. Using both engines , the ferryboat can make the crossing in 4hrs .How long would it take each engine , working alone ,to power the ferryboat across the channel

hero (hero):

1st engine x 2nd engine x - 6 \(\large\frac{x(x-6)}{x + (x - 6)} = 4\)

hero (hero):

Solve for x to get time it takes the first engine, then subtract 6 from x to get the second engine time.

OpenStudy (anonymous):

Let x be the time it takes the second engine to cross a channel alone. From the problem narrative, one can derive the following equation.\[\frac{1}{x+6}+\frac{1}{x}=\frac{1}{4} \]The solution for x is x = 6 hours x + 6 = 12 hours \[\frac{1}{12}+\frac{1}{6}=\frac{1}{4} \]

hero (hero):

That's very nice but solving for x using it my way also produces x = 12, x - 6 = 6 And then using the formula I provided you get: \(\frac{12(12-6)}{12 + 12 - 6}=4\) and \(\frac{72}{18}=4\)

OpenStudy (anonymous):

Yes, you let x be the first engine it seems, so the numeric values are reversed in the solutions. Don't remember doing "combined rate problems" in school. When first encountered on this site I memorized the method and have used the form of adding fractions, job/hour(s) since then.

hero (hero):

I used Satellite's \(\large\frac{M \times N}{M + N} = T\) method which I prefer.